OpenStudy (anonymous):

3y - x = 4 y + 2x = -1 choose the correct graph of given the system of equations/ find any solutions if any.

OpenStudy (callisto):

perhaps you should give us a graph?

OpenStudy (callisto):

3y - x = 4 => x = 3y-4 --(1) y + 2x = -1 => y+2x =-1--(2) put (1) into (2) y + 2(3y-4) = -1 7y = 7 y=1 put y=1 into (1) x= 3(1)-4 = -1 therefore x=-1 and y=1